Residual Income and Investment Decisions

Jarriot, Inc., presented two years of data for its Furniture Division and its Houseware Division.

Furniture Division:
        Year 1 Year 2
Sales $35,500,000 $38,400,000
Operating income 1,370,000 1,540,000
Average operating assets 8,590,000 8,590,000
Houseware Division:
        Year 1 Year 2
Sales $11,500,000 $12,700,000
Operating income 680,000 510,000
Average operating assets 5,800,000 5,800,000

At the end of Year 2, the manager of the Houseware Division is concerned about the division’s performance. As a result, he is considering the opportunity to invest in two independent projects. The first is called the Espresso-Pro; it is an in-home espresso maker that can brew regular coffee as well as make espresso and latte drinks. While the market for espresso drinkers is small initially, he believes this market can grow, especially around gift-giving occasions. The second is the Mini-Prep appliance that can be used to do small chopping and dicing chores that do not require a full-sized food processor. Without the investments, the division expects that Year 2 data will remain unchanged. The expected operating incomes and the outlay required for each investment are as follows:

Espresso-Pro Mini-Prep
Operating income $28,000 $15,300
Outlay 160,000 110,000

Jarriot’s corporate headquarters has made available up to $520,000 of capital for this division. Any funds not invested by the division will be retained by headquarters and invested to earn the company’s minimum required rate of return, 7 percent.

Required:

1. Compute the residual income for each of the opportunities. (Round to the nearest dollar.)

Espresso-Pro residual income $
Mini-Prep residual income $

2. Compute the divisional residual income for each of the following four alternatives: (Round to the nearest dollar.)

a. The Espresso-Pro is added.
$

b. The Mini-Prep is added.
$

c. Both investments are added.
$

d. Neither investment is made; the status quo is maintained.
$

Assuming that divisional managers are evaluated and rewarded on the basis of residual income, which alternative do you think the divisional manager will choose?

3. Assuming that management acts as you recommend in requirement 2, compute the change in profit (loss) from the divisional manager's investment decision.

   $

Was the correct decision made?
Answer #1
Part-1
Espresso-Pro ROI (28000/160000)= 17.5%
Mini-prep ROI (15300/110000)=13.91%
Part-2
a Espresso pro is added
(28000+510000)/(160000+5800000)=9.03%
b Mini prep is added
(15300+510000)/(110000+5800000)=8.89%
c Both are added
(15300+28000+510000)/(110000+160000+5800000)=9.12%
d Status quo (510000/5800000)=8.79%
ROI = Operating income/Operating assets
Espresso-Pro should be chosen as ROI is highest